OpenStudy (anonymous):

Given the equation y − 4 = three fourths(x + 8) in point-slope form, identify the equation of the same line in slope-intercept form. −three fourthsx + y = 10 3x − 4y = −40 y = three fourthsx + 12 y = three fourthsx + 10

OpenStudy (anonymous):

Distribute the 3/4 then isolate y

OpenStudy (anonymous):

what does distribute mean?

OpenStudy (anonymous):

multiply everything in the parentheses by 3/4

OpenStudy (anonymous):

so mutiply 3/4 with 8

OpenStudy (anonymous):

And x. so you get y-4=3/4x+6

OpenStudy (anonymous):

so i multiply both sides by 4

OpenStudy (anonymous):

i ment subtract.

OpenStudy (anonymous):

Yep. And you get y=3/4x+2 which is your answer.

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

You're welcome